OpenStudy (anonymous):

solve: (1.025)^12t = 3

OpenStudy (lgbasallote):

\[(1.025)^{12t} = 3\] first step take the log \[\Large \log_{1.025} (3) = 12t\] now divide by 12 to get t \[\large \frac{\log_{1.025} (3)}{12} = t\]

OpenStudy (anonymous):

if i wanted to get an answer with decimals. would i go log(3/1.025) / 12 ?

OpenStudy (lgbasallote):

no that's different

OpenStudy (lgbasallote):

just do that thing i said

OpenStudy (lgbasallote):

otherwise you'll have to go \[\frac{\log 3}{12\log 1.025} = t\]

OpenStudy (anonymous):

so i just leave it in the form you left it in as? log_1.025 (3) / 12 = t ... and i dont simplify it?

OpenStudy (lgbasallote):

you use a calculator to simplify of course

OpenStudy (anonymous):

im not sure how to plug it into my calculator

OpenStudy (lgbasallote):

do \[\log 3 \div \log 1.025\] what do you get

OpenStudy (anonymous):

44.492

OpenStudy (lgbasallote):

now divide by 12

OpenStudy (anonymous):

3.707

OpenStudy (lgbasallote):

\[t \approx 3.71\]`

OpenStudy (anonymous):

thanks!

OpenStudy (lgbasallote):

<tips imaginary hat>